If 10-year T-bonds have a yield of 6.2%, 10-year corporate bonds yield 10%, the maturity risk premium on all 10-year bonds is 1.3%, and corporate bonds have a 0.4% liquidity premium versus a zero liquidity premium for T-bonds, what is the default risk premium on the corporate bond?

Select the correct answer.

a. 2.65%
b. 2.90%
c. 3.40%
d. 2.40%
e. 3.15%
Answer #1

Solution:-

Risk free rate = 10 - 6.2 = 3.8 %

Default risk premium = Risk free rate -liquidity premium

                                        = 3.8 - 0.4

                                      = 3.40%

SO the answer is option (C) 3.40%